OpenStudy (anonymous):

@CrystalBurley Point C is graphed on a number line at –7. Point D is 13 units away from point C on the number line. What could be the coordinates of point D? Choose all answers that are correct. A. –20 B. 3 C. 6 D. 20

OpenStudy (anonymous):

a and d

OpenStudy (anonymous):

k i think it is a negative 20 so -20

OpenStudy (anonymous):

so a and b?

OpenStudy (anonymous):

rightt

OpenStudy (anonymous):

yea its right

OpenStudy (anonymous):

Which equation demonstrates the opposite of a sum property? A. (22 + 10) + 3 = (10 + 22) + 3 B. –5 + 13 = 13 + (–5) C. (12 + (–4)) + 14 = 12 + ((–4) + 14) D. –(–6 + 17) = 6 + (–17)